Two identical charges of 5nc repel each other when they are 5 cm apart from each other. find out the magnitude of force b/w them.


1
Expert's answer
2021-01-31T03:40:29-0500
"F=\dfrac{kq_1q_2}{r^2},""F=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot(5\cdot10^{-9}\ C)^2}{(0.05\ m)^2}=9.0\cdot10^{-5}\ N."